# ESPRESSO observations of the debris-accreting white dwarf WD 0141–675

### v1

@@ Abstract
WD 0141–675 was reported as the first astrometrically detected white dwarf planet host candidate as part of the third data release from Gaia, just to be later retracted via a news item on the Gaia web site1. We present time-resolved, high-resolution optical ESPRESSO spectroscopy of WD 0141–675. A radial velocity analysis of the Ca K absorption line reveals a tentative periodic signal of d. Phase-folding the ESPRESSO spectroscopy on this signal exhibits weak variability in the morphology of Ca K close to the core of the line. A violet-to-red ratio analysis of the Ca K line shows a periodic signal of d. The periods from both methods agree, within their uncertainties, with half the period of the astrometric planet candidate, however, both measurements are of low statistical significance. Nonetheless, our results imply possible solutions to the mass function within the planetary regime. And when combined with existing infrared photometry, which rules out a brown dwarf companion, yield a lower limit on the orbital inclination of . Our study demonstrates that ESPRESSO observations are well capable of detecting short-period (days to weeks) giant planets orbiting white dwarfs.
@@
Astrometry provides another method to identify planets orbiting nearby white dwarfs, and the potential of the Gaia mission in that respect has been thoroughly explored (Silvotti et al., 2011; Sanderson et al., 2022). Indeed, as part of its third data release (Gaia DR3; Gaia Collaboration et al. 2023a; Holl et al. 2023), the first astrometric planet candidate was announced: WD 0141–675.
@@ 2 WD 0141–675
WD 0141–675 is a nearby ( pc, Gaia Collaboration et al. 2023b) white dwarf with a hydrogen-dominated atmosphere (Hintzen & Jensen, 1979) that attracted little attention since its discovery, apart from the detection of weak Ca H&K absorption lines (Debes & Kilic, 2010), indicating ongoing accretion of planetary material. Interest in WD 0141–675 was sparked when an astrometric planet candidate with an orbital period of d and a mass of was announced (Gaia Collaboration et al., 2023a; Holl et al., 2023). However in 2023 May, the planet candidate was quietly retracted via a news item111https://www.cosmos.esa.int/web/gaia/dr3-known-issues, and classified as a false positive in the Gaia astrometric solutions.
@@ 3 Observations
We acquired high-resolution spectroscopy of WD 0141–675 with the ESPRESSO spectrograph at the VLT of the European Southern Observatory (ESO). We used the single-telescope configuration, slow read-out and binning (). We obtained a total of 20 spectra, each with an exposure time of 3144 s, throughout the period 2023 October 10 to 2023 December 20, details of the observations are provided in Table 3. The observations were designed to sample the putative orbital period of d, obtaining one spectrum on average once every two days. However, due to varying sky conditions, several nights were skipped, and the observations ended up being spread out over two months.
The ESPRESSO data were reduced with the Data Reduction Software222www.eso.org/sci/software/pipelines/espresso/ espresso-pipe-recipes.html (DRS) version 3.0.0. The DRS performs standard echelle spectrum reduction, including bias and dark subtraction, optimal order extraction, bad pixel correction, flat-fielding, deblazing, wavelength calibration, and order merging (for details see Pepe et al., 2021). Here we used the deblazed, order-merged spectra. We removed telluric lines using molecfit (Kausch et al., 2015; Smette et al., 2015). The signal-to-noise ratio (S/N) of the individual spectra, measured in between the Ca H&K absorption lines, ranges from 21 to 45 (Table 3).
@@ 4 Atmospheric analysis
We used the extremely high resolution of the ESPRESSO spectrum to place an upper limit on the rotational velocity, , corroborating earlier findings that most white dwarfs have rotation periods of many hours to several days (Hermes et al., 2017), too long to detect rotational broadening even in high-resolution spectroscopy (Koester et al., 1998; Berger et al., 2005).
@@ 5.1 Radial velocity analysis
We fitted a Gaussian profile to the Ca K absorption line to measure the radial velocity at the time of each exposure. The Lomb-Scargle (LS) periodogram computed from these radial velocities (left panel of Fig. 2) displays a strong peak at d, with a False Alarm Rate (FAR) of 99 per cent; a signal far too weak to be considered conclusive. Another, slightly weaker signal is detected at d. We pre-whitened the radial velocities with the d signal, and re-computed the LS periodogram, which caused the d signal to drop by per cent in power. Finally, upon further subtraction of the d signal, no signals remained in the LS.
@@
where is the systemic velocity, the amplitude, and was fixed to our selected period. The value of is that of the inferior conjunction, assuming the presence of a planet. The parameters of the optimal fit are listed on Table 1. The systemic velocity derived here, , is somewhat lower than the average of the radial velocities reported by Rogers et al. (2024), however, systematic uncertainties in the wavelength zero-point of ESPRESSO (our data) and MIKE and X-Shooter (Rogers et al., 2024) can account for this difference, in particular as the latter two instruments are not designed for stable, high-precision radial velocity studies. The amplitude determined from the ESPRESSO spectra, is well below the radial velocity precision of the MIKE data, (Rogers et al., 2024).
@@ 5.2 V/R ratio Analysis
We measured the violet-to-red (V/R) ratio for each ESPRESSO spectrum (e.g. Manser et al., 2019), which is the ratio of blue-shifted to red-shifted flux, centred on the air wavelengths of the Ca K line shifted by the systemic velocity (obtained in Section 5.1). The LS periodogram computed from the V/R ratios (Fig. 3, left panel) contains a strong signal at d. We note however, that the signal has a high FAR (99 per cent). The periodogram shows two additional signals at d and d.
The results of the V/R analysis are sensitive to the interval adopted in the integration of the V/R ratio, whereby wider intervals favoured the strength of the shorter periods, while narrow intervals increased the strength of the day-period. Our final choice was an interval of to around the core of the line, which yields all three signals.
Subtracting the 16 d signal decreases the power of the d signal by 50 per cent and effectively removes the d signal, analogous to the radial velocity LS (Section 5.1), suggesting these shorter periods as being aliases introduced by the temporal sampling of the ESPRESSO observations.
We phase-folded the data on the d-period and fitted a sine function to the data. The best fit is shown in the right panel of Fig. 3, and the corresponding parameters are listed on Table 2. We note that, as in the RV measurement discussed above, the periodic signal detected in the V/R data is roughly half of that reported for the retracted planet candidate ( d).
To visualise the V/R variability, we first computed the average spectrum of all the ESPRESSO observations, subtracted that average from each individual spectrum, and then phase-folded the data on the d period, with the result displayed as a phase-trailed spectrogram in Fig. 4. Two tentative sine wave patterns are present in the residual spectrogram (bottom panel of Fig. 4).
@@ 6 Discussion
In Section 5.1, we tentatively detected a radial velocity variation in the Ca K line with 0.9 d and an amplitude of . Taking these parameters at face value and interpreting them as the circular orbital motion of the white dwarf resulting from a close-in planet leads to the following mass function for the putative companion:
where is the gravitational constant and the mass of the WD as reported by Rogers et al. (2024). This mass function constrains the possible values for the inclination () and the mass of the companion (), as illustrated in Fig. 5. For an orbital inclination the companion would have a mass in the brown dwarf regime, whereas allows only for a planetary mass companion. The mass reported by Gaia, , yields an orbital inclination of , which is in stark disagreement with the inclination of inferred from Monte Carlo resampling by Gaia Collaboration et al. (2023b). However, we remind the reader that the astrometric solution for WD 0141–675 has been retracted by the Gaia team. We also note that while the available Spitzer and WISE photometry rules out a brown dwarf companion (see section 3.3 of Rogers et al. 2024 for a detailed discussion), a more stringent photometric limit on the existence of a planetary companion could be derived from deep JWST infrared observations333The approved JWST program GO 3652 intended to obtain these observations, but was withdrawn in the wake of the retraction of WD 0141–675. We argue that a better constraint on the nature of a possible companion to WD 0141–675 remains of interest..
@@ 7 Conclusions
WD 0141–675 was reported as hosting the first astrometrically detected planet candidate (Gaia Collaboration et al., 2023a; Holl et al., 2023), but was later retracted as a false positive. We have obtained time-resolved high-resolution spectroscopy of WD 0141–675 with ESPRESSO, an instrument which allows highly accurate velocity shift measurements (e.g. Pasquini et al., 2023). Adopting the atmospheric parameters from the photometric solution of Rogers et al. (2024), we obtained a Ca abundance of dex, in good agreement with the values reported by these authors. The narrow width of the Ca K lines yielded an upper limit on the rotational broadening of . This is suggestive of a slow rotator, which is a common result in spectroscopic determinations of rotating rates in white dwarfs (Berger et al., 2005).
Analysing the radial velocities and the V/R ratio of the Ca K line both result in a tentative detection of variability with a period of d. Phase-folding the ESPRESSO spectroscopy on such signal exhibits weak periodical variability in the morphology of Ca K near its core. Fitting the radial velocities with a circular orbit results in an amplitude of . Interpreting this variation as the (circular) reflex motion of the white dwarf implies that possible solutions to the mass function exist within the planetary mass regime. Therefore, when discarding a brown dwarf companion from existing infrared photometry (Rogers et al., 2024) a lower limit on the orbital inclination of is established.
@@ Acknowledgements
We thank James Munday for useful discussions on radial velocity methods. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Framework Programme (grant agreement no. 101020057). This research was in part funded by the UKRI (Grant EP/X027562/1) Research based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme(s) 112.2604.001.
@@ Appendix A Log of the observations
Date | Time | S/N |
2023 Oct 20 | 03:13:49 | 34 |
2023 Oct 21 | 02:16:07 | 36 |
2023 Oct 25 | 06:03:14 | 34 |
2023 Oct 29 | 02:32:04 | 32 |
2023 Nov 05 | 03:57:32 | 33 |
2023 Nov 11 | 01:32:23 | 34 |
2023 Nov 13 | 01:54:27 | 29 |
2023 Nov 14 | 01:52:00 | 24 |
2023 Nov 19 | 01:41:56 | 24 |
2023 Nov 26 | 00:39:32 | 45 |
2023 Dec 05 | 00:30:18 | 37 |
2023 Dec 06 | 00:46:26 | 30 |
2023 Dec 07 | 01:55:34 | 33 |
2023 Dec 08 | 00:52:07 | 28 |
2023 Dec 09 | 01:47:59 | 21 |
2023 Dec 10 | 03:21:44 | 24 |
2023 Dec 11 | 03:30:44 | 32 |
2023 Dec 12 | 01:11:57 | 28 |
2023 Dec 16 | 01:28:59 | 30 |
2023 Dec 20 | 02:23:25 | 31 |

### v2

@@ Abstract
WD 0141–675 was reported as the first astrometrically detected white dwarf planet host candidate as part of the third data release from Gaia, just to be later retracted via a news item on the Gaia web site1. We present time-resolved, high-resolution optical ESPRESSO spectroscopy of WD 0141–675. A radial velocity analysis of the Ca K absorption line reveals a tentative periodic signal of d and an amplitude modulation of km s-1. Phase-folding the ESPRESSO spectroscopy on this signal exhibits weak variability in the morphology of Ca K close to the core of the line. A violet-to-red ratio analysis of the Ca K line shows a periodic signal of d. The periods from both methods agree, within their uncertainties, with half the period of the astrometric planet candidate, however, both measurements are of low statistical significance. Nonetheless, our results imply possible solutions to the mass function within the planetary regime. When combined with existing infrared photometry, which rules out a brown dwarf companion, yield a lower limit on the orbital inclination of . Our study demonstrates that ESPRESSO observations are well capable of detecting short-period (days to weeks) giant planets orbiting white dwarfs.
@@
Astrometry provides another method to identify planets orbiting nearby white dwarfs, and the potential of the Gaia mission in that respect has been thoroughly explored (Silvotti et al., 2011; Sanderson et al., 2022). Indeed, as part of its third data release (Gaia DR3; Gaia Collaboration et al. 2023a; Holl et al. 2023), the first astrometric white dwarf planet candidate was announced: WD 0141–675.
@@ 2 WD 0141–675
WD 0141–675 is a nearby ( pc, Gaia Collaboration et al. 2023b) white dwarf with a hydrogen-dominated atmosphere (Hintzen & Jensen, 1979), with weak Ca H&K absorption lines (Debes & Kilic, 2010), indicating ongoing accretion of planetary material. Interest in WD 0141–675 was recently sparked when an astrometric planet candidate with an orbital period of d and a mass of was announced (Gaia Collaboration et al., 2023a; Holl et al., 2023). However in 2023 May, the planet candidate was quietly retracted via a news item111https://www.cosmos.esa.int/web/gaia/dr3-known-issues, and classified as a false positive in the Gaia astrometric solutions.
@@ 3 Observations
We acquired high-resolution spectroscopy of WD 0141–675 with the ESPRESSO spectrograph at the VLT of the European Southern Observatory (ESO). We used the single-telescope configuration, slow read-out and binning (). We obtained a total of 20 spectra, each with an exposure time of 3144 s, throughout the period 2023 October 10 to 2023 December 20; further details of the observations are provided in Table 3. The observations were designed to sample the putative orbital period of d, obtaining one spectrum on average once every two days. However, due to varying sky conditions, several nights were skipped, and the observations ended up being spread out over two months.
The ESPRESSO data were reduced with the Data Reduction Software222www.eso.org/sci/software/pipelines/espresso/espresso-pipe-recipes.html (DRS) version 3.0.0. The DRS performs standard echelle spectrum reduction, including bias and dark subtraction, optimal order extraction, bad pixel correction, flat-fielding, deblazing, wavelength calibration, and order merging (for details see Pepe et al., 2021). Here we used the deblazed, order-merged spectra. We removed telluric lines using molecfit (Kausch et al., 2015; Smette et al., 2015). The signal-to-noise ratio (S/N) of the individual spectra, measured around the Ca K absorption line, ranges from 14 to 25 (Table 3).
@@ 4 Atmospheric analysis
We used the extremely high resolution of the ESPRESSO spectrum to place an upper limit on the rotational velocity, (See Figure 1), corroborating earlier findings that most white dwarfs have rotation periods of many hours to several days (Hermes et al., 2017), too long to detect rotational broadening even in high-resolution spectroscopy (Koester et al., 1998; Berger et al., 2005).
@@ 5.1 Radial velocity analysis
We fitted a Gaussian profile to the Ca K absorption line to measure the radial velocity (RV) at the time of each exposure. We then computed the Lomb-Scargle periodogram from the astropy package (Astropy Collaboration et al., 2022), and calculated the False Alarm Rate (FAR) using the option based on the formulation given by Baluev (2008).
The RVs periodogram (left panel of Fig. 2) displays a strong peak at d, roughly consistent with half the period of the astrometric planet candidate (=33.65, Gaia Collaboration et al., 2023a; Holl et al., 2023), although with a FAR of 99 per cent. Another slightly weaker signal is also detected at d.
Besides having such high FAR, each of the found signals present additional reasons for caution: The 15 d signal overlaps with a peak rising from the spectral window function, hinting at a possible link to the temporal sampling; and the 10 d signal coincides with a third of the lunar cycle. In Section 6.1, we explore in more detail these systematics, and argue that the d signal stands as our best orbital period candidate.
Upon pre-whitening the 10 d period, the 15 d signal became stronger by 30 per cent. And finally, pre-whitening both periods did not uncover new signals in the LS .
@@
where is the systemic velocity, the amplitude, and is the orbital period. The value of is that of the inferior conjunction, assuming the presence of a planet. The parameters of the optimal fit are listed on Table 1.
The systemic velocity derived here, , is somewhat lower than the average of the radial velocities reported by Rogers et al. (2024), however, systematic uncertainties in the wavelength zero-point of ESPRESSO (our data), MIKE and X-Shooter (Rogers et al., 2024) can account for this difference. The latter two instruments are not designed for stable, high-precision radial velocity studies. Indeed, the amplitude determined from the ESPRESSO spectra, is well below the radial velocity precision of the MIKE data, (Rogers et al., 2024).
@@ 5.2 V/R ratio Analysis
We measured the violet-to-red (V/R) ratio for each ESPRESSO spectrum (e.g. Manser et al., 2019), which is the ratio of blue-shifted to red-shifted flux, centred on the air wavelengths of the Ca K line shifted by the systemic velocity (obtained in Section 5.1). The LS periodogram computed from the V/R ratios (Fig. 3, left panel) contains a strong signal at d, which agrees within the uncertainties with the period found in our RV analysis in Section 5.1. We note however, that the signal has a high FAR (99 per cent). The periodogram shows two additional comparable signals at d and d.
In contrast to the RV analysis in Section 5.1, pre-whitening the 10 d signal decreased the power of both the d and the d signals. Further pre-whitening of the d period did not reveal new signals in the LS.
The results of the V/R analysis are sensitive to the interval adopted in the integration of the V/R ratio, whereby wider intervals favoured the strength of the shorter periods, while narrow intervals increased the strength of the d period. Our final choice was an interval of to around the core of the line, which yields all three signals with a comparable strength.
We phase-folded the data with the d-period and fitted a sine function to the data. The best fit is shown in the right panel of Fig. 3, and the corresponding parameters are listed on Table 2. As in the RV measurement discussed above, the periodic signal detected in the V/R data is consistent with half of that reported for the retracted planet candidate ( d). On the other hand, it is worth noting that the amplitude modulation and systemic offset are dimensionless parameters in this analysis, and are therefore not comparable to the parameters obtained in the RV study from Section 5.1.
To visualise the variability of the spectra along the orbit, we first computed the average spectrum of all the ESPRESSO observations, subtracted that average from each individual spectrum, and then phase-folded the data on the d period, with the result displayed as a phase-trailed spectrogram in Fig. 4. Two tentative sine wave patterns are present in the residual spectrogram (bottom panel of Fig. 4), but a spurious origin cannot be discarded without more observations.
@@ 6 Discussion
6.1 Systematics Test
We explored two possible systematic origins for the periodic signals arising in the radial velocity and V/R measurements: The temporal sampling of the ESPRESSO spectra and the lunar cycle. For this purpose we produced 10 000 bootstrap samples of our measurements for which we computed LS periodograms, recording the period with the strongest power in each sample. We then repeated the experiment imposing constraints on the angular separation from the Moon (), the fraction lunar illumination (FLI, as well as on the S/N () around Ca K. We show the histograms of the most frequent strongest periods found in the bootstrap samples in Fig. 5. When using the complete set of spectra the same signals dominating the LS periodograms in Sections 5.1 and 5.2 are those appearing most frequently in the bootstrap histograms. The only exception is the d signal in the radial velocity histogram, which is consistent to the harmonic in frequency of the 16 d signal. These results suggest that the temporal-sampling is not the cause of the detected signals.
When imposing the lunar and S/N constraints on the radial velocities (top panel of Fig. 5), the overall structure remains mostly unaffected. However, in the case of the V/R measurements the signal at d increases its occurrence rate at the expense of a decrease in the occurrence of the 10 and 12 d signals.
This marginally hints at sky background contamination as a possible origin for these shorter signals. For this reason, we argue that the signal at d is the best orbital period candidate; but we still issue caution, as further observations are needed to sustain such assessment.
6.2 Implications on the Mass Function
In Section 5.1, we tentatively detected a radial velocity variation in the Ca K line with 0.9 d and an amplitude of . Furthermore, a consistent period within the errors of d was also found in the V/R analysis (Section 5.2). This period agrees within uncertainties with that of the astrometric period candidate ( d, Gaia Collaboration et al., 2023a; Holl et al., 2023)
Taking the RV parameters at face value and interpreting them as the circular orbital motion of the white dwarf resulting from a close-in planet leads to the following mass function for the putative companion:
where is the gravitational constant and the mass of the white dwarf as reported by Rogers et al. (2024). This mass function constrains the possible values for the inclination () and the mass of the companion (), as illustrated in Fig. 6. For an orbital inclination the companion would have a mass in the brown dwarf regime, whereas allows only for a planetary mass companion. The companion mass reported by Gaia, , yields an orbital inclination of in our analysis, which is in stark disagreement with the inclination of inferred from Monte Carlo resampling by Gaia Collaboration et al. (2023b). However, we remind the reader that the astrometric solution for WD 0141–675 has been retracted by the Gaia team. We also note that while the available Spitzer and WISE photometry rules out a brown dwarf companion (see section 3.3 of Rogers et al. 2024 for a detailed discussion), a more stringent photometric limit on the existence of a planetary companion could be derived from deep JWST infrared observations333The approved JWST program GO 3652 intended to obtain these observations, but was withdrawn in the wake of the retraction of WD 0141–675. We argue that a better constraint on the nature of a possible companion to WD 0141–675 remains of interest..
@@ 7 Conclusions
WD 0141–675 was reported as hosting the first astrometrically detected white dwarf planet candidate (Gaia Collaboration et al., 2023a; Holl et al., 2023), but was later retracted as a false positive. We have obtained time-resolved high-resolution spectroscopy of WD 0141–675 with ESPRESSO, an instrument which allows highly accurate velocity shift measurements (e.g. Pasquini et al., 2023). Adopting the atmospheric parameters from the photometric solution of Rogers et al. (2024), we obtained a Ca abundance of dex, in good agreement with the values reported by these authors. The narrow width of the Ca K lines yielded an upper limit on the rotational broadening of . This is suggestive of a slow rotator, which is a common result in spectroscopic determinations of rotating rates in white dwarfs (Berger et al., 2005).
Analysing the radial velocities and the V/R ratio of the Ca K line both result in a tentative detection of variability with a period of d, which cannot be attributed to the temporal sampling. Phase-folding the ESPRESSO spectroscopy on such signal exhibits weak periodical variability in the morphology of Ca K near its core. Fitting the radial velocities with a circular orbit results in an amplitude of . Interpreting this variation as the (circular) reflex motion of the white dwarf implies that possible solutions to the mass function exist within the planetary mass regime. Therefore, when discarding a brown dwarf companion from existing infrared photometry (Rogers et al., 2024) a lower limit on the orbital inclination of is established.
@@ Acknowledgements
We thank James Munday for useful discussions on radial velocity methods. We would also like to thank the anonymous referee, whose comments helped improve the content of this article. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Framework Programme (grant agreement no. 101020057). This research was in part funded by the UKRI (Grant EP/X027562/1) Research based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme(s) 112.2604.001.
@@ Appendix A Log of the observations
Date | Time | MJD | Lunar | FLI | S/N |
Dist. | |||||
2023-10-20 | 03:13:49 | 60237.134604 | 74 | 0.3 | 20.9 |
2023-10-21 | 02:16:07 | 60238.094533 | 69 | 0.4 | 21.5 |
2023-10-25 | 06:03:14 | 60242.252246 | 64 | 0.8 | 20.1 |
2023-10-29 | 02:32:04 | 60246.105605 | 84 | 1.0 | 20.3 |
2023-11-05 | 03:57:32 | 60253.164957 | 118 | 0.5 | 20.2 |
2023-11-11 | 01:32:23 | 60259.064163 | 105 | 0.1 | 19.9 |
2023-11-13 | 01:54:27 | 60261.079487 | 94 | 0.0 | 17.5 |
2023-11-14 | 01:52:00 | 60262.077783 | 88 | 0.0 | 14.2 |
2023-11-19 | 01:41:56 | 60267.070795 | 63 | 0.3 | 14.1 |
2023-11-26 | 00:39:32 | 60274.027459 | 87 | 1.0 | 25.8 |
2023-12-05 | 00:30:18 | 60283.021052 | 117 | 0.5 | 22.2 |
2023-12-06 | 00:46:26 | 60284.032248 | 115 | 0.4 | 18.1 |
2023-12-07 | 01:55:34 | 60285.080263 | 111 | 0.3 | 19.5 |
2023-12-08 | 00:52:07 | 60286.036197 | 107 | 0.2 | 16.3 |
2023-12-09 | 01:47:59 | 60287.074998 | 102 | 0.2 | 12.0 |
2023-12-10 | 03:21:44 | 60288.140102 | 96 | 0.1 | 13.9 |
2023-12-11 | 03:30:44 | 60289.146346 | 89 | 0.0 | 19.3 |
2023-12-12 | 01:11:57 | 60290.049968 | 84 | 0.0 | 16.8 |
2023-12-16 | 01:28:59 | 60294.061797 | 64 | 0.1 | 18.0 |
2023-12-20 | 02:23:25 | 60298.099596 | 69 | 0.5 | 19.0 |